ANALYSIS OF SINGLET AND TRIPLET STRUCTURE IN THE HIGH-RESOLUTION SPECTRUM OF ZIRCONIUM MONOCARBIDE

ANTHONY J. MERER and JAMES R.D. PEERS^{a}, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, BC, Canada V6T 1Z1; \underline{SCOTT J. RIXON}, Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC, Canada V6T 1Z1.

ZrC is readily produced in supersonic expansion via reaction of 1% methane (in helium) with laser-ablated zirconium metal. Laser-induced fluorescence excitation and dispersed emission spectra have been recorded for bands in the region 630 − 520 nm (16000 − 19000 cm^{−1}). The four lowest electronic states lie below 2500 cm^{−1} and appear to represent all possible arrangements of two electrons in the two nearly degenerate 11σ and 12σ orbitals (both formed from Zr 5sσ + C 2pσ); the evidence is that the small spin-spin interaction of the X^{3}Σ^{+} ground state (λ_{0} = 0.5142 cm^{−1}, r_{0} = 1.807 Å for ^{90}Zr^{12}C) is consistent with a σσ′ configuration, while the anomalous ^{12}C/^{13}C isotope shifts and vibrational intervals of the remaining three states show that they all have the same (^{1}Σ^{+}) symmetry. In particular, the tightly bound a^{1}Σ^{+} state (T_{0} = 187.83 cm^{−1}, r_{0} = 1.739 Å) arises from the 11σ^{2} closed shell configuration. Moreover, this molecular orbital scheme is consistent with the confused level structure above 16000 cm^{−1}. At least four close-lying Π states (two singlets and two triplets) are observed and can be explained by the promotion of either σ electron to a π orbital; these exhibit both isoconfigurational (^{1}Π/^{3}Π_{1}) and interconfigurational (^{1}Π/^{1}Π or ^{3}Π/^{3}Π) interactions. Details of the rotational and 91Zr (I = 5/2) hyperfine structures will be presented.

aPresent address: Creo Inc., 3755 Willingdon Avenue, Burnaby, BC, Canada V5G 3H3